R2, T(21,02) = (3.01 + 2.62,5x1 + 3x3) Find the formula for Problem 4. Let T : R2 T-131, 12)
Answer #1

Ques; let T3Bbe T61/25) = (3472x2,5x+379) Let x, y ER² and Care Then, x+ cy: (21,) + c (or ot) = + Go X27C42) So. T(*+ cy) =And T(0,1) = (0+2, 0+3) - (2,3) fet d = {(1,0), (01) be the standard ordered basis of B? Then, it is 2 7 (50,5- ] chains ((79> T( 0, Ms) = *(-3,5) + ? (2,-3) » 7 (3,2/2) = (-3x1 +242 7581-342) (Hence. The result)